Solve the system of linear equations using any method.
2x + y = 5 and 4x + 3y = 11
does this have?
The system has infinitely many solutions.
The system has one and only one solution (2, 1).
The system has no solutions.
none of the above?

Answer by ewatrrr(24785)

 
Hi,
Solve the system of linear equations using any method.
2x + y = 5 OR y = -2x+5
4x + 3y = 11
4x + 3(-2x+5) = 11
-2x = -4
x = 2 and y = 1 (-2*2 + 5 = 1)
The system has one and only one solution (2, 1).